# SynSUM – Synthetic Benchmark with Structured and Unstructured Medical Records

### v2

@@ Abstract
We present the SynSUM benchmark, a synthetic dataset linking unstructured clinical notes to structured background variables. The dataset consists of $10{,}000$ artificial patient records containing tabular variables (like symptoms, diagnoses and underlying conditions) and related notes describing the fictional patient encounter in the domain of respiratory diseases. The tabular portion of the data is generated through a Bayesian network, where both the causal structure between the variables and the conditional probabilities are proposed by an expert based on domain knowledge. We then prompt a large language model (GPT-4o) to generate a clinical note related to this patient encounter, describing the patient symptoms and additional context. We conduct both an expert evaluation study to assess the quality of the generated notes, as well as running some simple predictor models on both the tabular and text portions of the dataset, forming a baseline for further research. The SynSUM dataset is primarily designed to facilitate research on clinical information extraction in the presence of tabular background variables, which can be linked through domain knowledge to concepts of interest to be extracted from the text - the symptoms, in the case of SynSUM. Secondary uses include research on the automation of clinical reasoning over both tabular data and text, causal effect estimation in the presence of tabular and/or textual confounders, and multi-modal synthetic data generation.
Code & Datasethttps://github.com/prabaey/SynSUM
1 Introduction
@@
Electronic health records (EHRs) are a gold mine of information, containing a mix of structured tabular variables (medication, diagnosis codes, lab results…) and free unstructured text (detailed clinical notes from physicians, nurses…) (Ford et al. 2016). These EHRs form a valuable basis for training clinical decision support systems, (partially) automating essential processes in the clinical world, such as diagnosis, writing treatment plans, and more (Peiffer-Smadja et al. 2020; Mujtaba et al. 2019; Rasmy et al. 2021; Li et al. 2020; Xu et al. 2019). While large language models can help leverage the potential of the unstructured text portion of the EHR (Zhang et al. 2020; Liu et al. 2022; Huang, Altosaar, and Ranganath 2019; Lehman and Johnson 2023; Singhal et al. 2023; Labrak et al. 2024), these black box systems lack interpretability (Quinn et al. 2022; Zhao et al. 2024; Tian et al. 2024). In high-risk clinical applications, it can be argued that one should prefer more robust and transparent systems built on simpler, feature-based models, like regression models, decision trees, or Bayesian networks (Rudin 2019; Sanchez et al. 2022; Lundberg et al. 2020). However, such models cannot directly deal with unstructured text and require tabular features as an input. For this reason, automated clinical information extraction (CIE) (Ford et al. 2016; Wang et al. 2018; Hahn and Oleynik 2020) is an essential tool for building large structured datasets that can serve as training data for such systems.
@@
However, CIE remains a challenging task due to the complex nature of clinical notes. These often leave out important contextual details which an automated system would need in order to correctly extract concepts from the text. Existing systems do not fully exploit the available medical domain knowledge to fill in this gap. We propose that CIE could benefit from leveraging two additional sources of information, apart from the unstructured text itself. On the one hand, a range of tabular features are already encoded in the EHR. These contain information related to a particular patient visit (e.g. partially encoded symptoms or diagnosis codes), as well as information on the medical history of the patient. On the other hand, we can connect this encoded background information with the concepts we are trying to extract from the text, using a Bayesian network (BN) that represents medical domain knowledge. A visual example that helps illustrate this idea is shown in Figure 2 in Appendix A.
@@
To investigate and implement the described research idea, we need a clinical dataset which (i) contains a mix of tabular data and unstructured text, where (ii) the tabular data and the concepts we aim to extract from the text can be linked through domain knowledge. While open-source datasets like MIMIC-III (Johnson et al. 2016) and MIMIC-IV (Johnson et al. 2023) contain this mix, they are not a perfect fit. First, the area of intensive care in which the data was collected is very extensive, making it hard to isolate a specific small-scale use-case for which the domain knowledge could be listed. Second, the portion of the dataset which is encoded into tabular features is often driven by billing needs, rather than completeness or accuracy, and does not contain any encoded symptoms, which are concepts that could be interesting to extract from the text for application in clinical decision support systems. Third, the link between the tabular features and the concepts mentioned in the text might be inconsistent due to system design or human errors (Kwon et al. 2024). Finally, the EHRs in MIMIC being time series adds another layer of complexity. Other existing datasets linking unstructured clinical text to structured features include BioDEX (D’Oosterlinck et al. 2023), a large set of papers describing adverse drug events, as well as TCGA-Reports (Kefeli and Tatonetti 2024), a set of cancer pathology reports, both accompanied by tabular patient descriptors and extracted biological features. However, in both cases, it is not trivial to devise a BN representing the relevant expert knowledge, partially because that knowledge is not fully understood yet.
@@
In this work, we build a synthetic yet sufficiently realistic dataset that addresses some of these shortcomings, enabling research on incorporating domain knowledge for improved CIE in the presence of tabular variables. Our dataset, called SynSUM (Synthetic Structured and Unstructured Medical records) is a self-contained set of synthetic EHRs in a primary care setting, fulfilling the following requirements:
@@
SynSUM is constructed in the domain of respiratory diseases, simulating patient visits to a primary care doctor. We mimic the scenario where the doctor notes down the patient’s symptoms in a clinical note, along with some additional context, and stores this in the EHR together with the encoded diagnosis, as well as the encoded symptoms. Additionally, the EHR stores tabular background information on the underlying health conditions of the patient.
Section 2 describes how we generated 10,000 fictional patient encounters by sampling from an expert-defined BN and prompting a large language model. Section 3 describes the expert evaluation conducted by a panel of five primary care physicians on a subset of the generated notes. Section 4 then runs some simple symptom predictors on both the tabular and textual portions of our dataset, which form a baseline to be exceeded by future research. Finally, Section 5 discusses the potential use of our dataset in various areas of research (including but not limited to CIE), while also highlighting its limitations.
2 Data generation
Figure 1: Overview of the full data generating process for the SynSUM dataset. First, the tabular portion of the synthetic patient record is sampled from a Bayesian network, where both the structure and the conditional probability distributions were defined by an expert. Afterwards, we construct a prompt describing the symptoms experienced by the patient, as well as their underlying health conditions (but no diagnoses). We ask the GPT-4o large language model to generate a clinical note describing this patient encounter. Finally, we ask to generate a more challenging compact version of the note, mimicking the complexity of real clinical notes by prompting the use of abbreviations and shortcuts. We generate $10{,}000$ of these synthetic patient records in total.
Our general methodology for generating the artificial patient records is shown in Figure 1. We now zoom in on the two major parts of this data generating process. First, Section 2.1 describes how we generated the structured tabular variables through an expert-defined Bayesian network (BN). Then, Section 2.2 dives into the clinical note generation by the large language model.
@@ 2.1 Modeling structured tabular variables with a Bayesian network
Causal structure
@@
We asked an expert to define a Directed Acyclic Graph (DAG) which (partially) models the domain of respiratory diseases in primary care, shown in Figure 1. In this DAG, a directed arrow between two variables models a causal relation between them. Central to the model are the diagnoses of pneumonia and common cold, which may give rise to five symptoms (dyspnea, cough, pain, fever and nasal). The expert also modeled some relevant underlying conditions which may render a patient more predisposed to certain diagnoses or symptoms: asthma, smoking, COPD and hay fever. Based on the symptoms experienced by a patient, a primary care doctor decides whether to prescribe antibiotics or not. The presence and severity of the symptoms, as well as the prescription of antibiotics as a treatment, influence the outcome, which is the total number of days (# days) that the patient eventually stays home as a result of illness. Finally, there are some non-clinical variables which exert an external influence on the diagnoses, treatment and outcome (season, policy and self-employed). Table 3 in Appendix B summarizes all variables and their meaning, as well as their possible values. While this model does not completely describe the real world, we do believe it to be sufficiently realistic for the purpose of generating a useful artificial dataset.
Probability distribution
@@
We turn the DAG into a BN by defining a joint distribution, which factorizes into $16$ conditional distributions, one for each variable. We use four different approaches to define these conditionals, after which we can sample synthetic patients from the BN in a top-down fashion.
For each of the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever, policy and self-employed, we ask our expert to define a conditional probability table (CPT). When a variable has many parents, it becomes infeasible to manually fill in the CPT in a clinically meaningful way. For this reason, we use a Noisy-OR distribution for the symptoms dyspnea, cough, pain and nasal, which assumes an independent causal mechanism behind the activation of a symptom through any of its parents. We model the prescription of antibiotics through a logistic regression model, mimicking the way the clinician’s suspicion of pneumonia (which needs to be treated with antibiotics) rises when a higher number of symptoms is present in the patient, with some symptoms weighing more than others, eventually exceeding a threshold. These weights are defined by the expert and validated with a set of test cases. In a similar fashion, we model the # days the patient stays home as a result of their complaints with a Poisson regression model. These weights are learned through maximum likelihood estimation over a set of train cases provided by the expert. The full specification of our BN can be found in Appendix B.

### v3

# SimSUM – Simulated Benchmark with Structured and Unstructured Medical Records

@@ Abstract
Background: Clinical information extraction, which involves structuring clinical concepts from unstructured medical text, remains a challenging problem that could benefit from the inclusion of tabular background information available in electronic health records. Existing open-source datasets lack explicit links between structured features and clinical concepts in the text, motivating the need for a new research dataset.
Methods: We introduce SimSUM [^1] An earlier version of this dataset was published under the name SynSUM. It has since been renamed to SimSUM to avoid confusion with synthetic data generated from real data, and to emphasize the simulated nature of the dataset., a benchmark dataset of 10,000 simulated patient records that link unstructured clinical notes with structured background variables. Each record simulates a patient encounter in the domain of respiratory diseases and includes tabular data (e.g., symptoms, diagnoses, underlying conditions) generated from a Bayesian network whose structure and parameters are defined by domain experts. A large language model (GPT-4o) is prompted to generate a clinical note describing the encounter, including symptoms and relevant context. These notes are annotated with span-level symptom mentions. We conduct an expert evaluation to assess note quality and run baseline predictive models on both the tabular and textual data.
Conclusion: The SimSUM dataset is primarily designed to support research on clinical information extraction in the presence of tabular background variables, which can be linked through domain knowledge to concepts of interest to be extracted from the text—namely, symptoms in the case of SimSUM. Secondary uses include research on the automation of clinical reasoning over both tabular data and text, causal effect estimation in the presence of tabular and/or textual confounders, and multi-modal synthetic data generation. SimSUM is not intended for training clinical decision support systems or production-grade models, but rather to facilitate reproducible research in a simplified and controlled setting. The dataset is available at https://github.com/prabaey/SimSUM.
keywords:
Bayesian network, clinical notes, clinical information extraction, electronic health records, large language model, tabular data, simulated benchmark.
1 Background
@@
Electronic health records (EHRs) are a gold mine of information, containing a mix of structured tabular variables (medication, diagnosis codes, lab results…) and free unstructured text (detailed clinical notes from physicians, nurses…) [4]. These EHRs form a valuable basis for training clinical decision support systems, (partially) automating essential processes in the clinical world, such as diagnosis, writing treatment plans, and more [22, 20, 25, 17, 32]. While large language models can help leverage the potential of the unstructured text portion of the EHR [33, 18, 8, 16, 29, 14], these black box systems lack interpretability [23, 34, 30]. In high-risk clinical applications, it can be argued that one should prefer more robust and transparent systems built on simpler, feature-based models, like regression models, decision trees, or Bayesian networks [27, 28, 19]. However, such models cannot directly deal with unstructured text and require tabular features as an input. For this reason, automated clinical information extraction (CIE) [4, 31, 6] is an essential tool for building large structured datasets that can serve as training data for such systems.
@@
Clinical information extraction remains a challenging task due to the complex nature of clinical notes. These often leave out important contextual details which an automated system would need in order to correctly extract concepts from the text. Existing systems do not fully exploit the available medical domain knowledge to fill in this gap. We propose that CIE could benefit from leveraging two additional sources of information, apart from the unstructured text itself. On the one hand, a range of tabular features are already encoded in the EHR. These contain information related to a particular patient visit (e.g. partially encoded symptoms or diagnosis codes), as well as information on the medical history of the patient. On the other hand, we can connect this encoded background information with the concepts we are trying to extract from the text, using a Bayesian network (BN) that represents medical domain knowledge. A visual example that helps illustrate this idea is shown in Fig. 1.
@@
To investigate and implement this idea, we need a clinical dataset which (i) contains a mix of tabular data and unstructured text, where (ii) the tabular data and the concepts we aim to extract from the text can be linked through domain knowledge. While open-source datasets like MIMIC-III [10] and MIMIC-IV [9] contain this mix, they are not a perfect fit. First, the area of intensive care in which the data was collected is very extensive, making it hard to isolate a specific use-case for which the domain knowledge could be listed. Second, the portion of the dataset which is encoded into tabular features is often driven by billing needs, rather than completeness or accuracy, and does not contain any encoded symptoms, which are concepts that could be interesting to extract from the text for application in clinical decision support systems. Third, the link between the tabular features and the concepts mentioned in the text might be inconsistent due to system design or human errors [13]. Finally, the EHRs in MIMIC being time series adds another layer of complexity. Other existing datasets linking unstructured clinical text to structured features include BioDEX [3], a large set of papers describing adverse drug events, as well as TCGA-Reports [11], a set of cancer pathology reports, both accompanied by tabular patient descriptors and extracted biological features. However, in both cases, it is not trivial to devise a BN representing the relevant expert knowledge, partially because that knowledge is not fully understood yet.
@@
In this work, we simulate a sufficiently realistic dataset that addresses some of these shortcomings, enabling research on incorporating domain knowledge for improved CIE in the presence of tabular variables. Our dataset, called SimSUM (Simulated Structured and Unstructured Medical records) is a self-contained set of artificial EHRs in a primary care setting, fulfilling the following requirements:
@@
SimSUM is constructed in the domain of respiratory diseases, simulating patient visits to a primary care doctor. We mimic the scenario where the doctor notes down the patient’s symptoms in a clinical note, along with some additional context, and stores this in the EHR together with the encoded diagnosis, as well as the encoded symptoms. Additionally, the EHR stores tabular background information on the underlying health conditions of the patient.
It is important to emphasize that the SimSUM dataset is not intended for training clinical decision support systems or deploying predictive models in real-world healthcare. Instead, it provides a controlled research environment for developing and prototyping methods. Although designed to be domain-relevant, SimSUM is entirely synthetic: tabular data is generated via an expert-defined Bayesian network, and clinical notes are produced by a large language model. It does not reflect the full variability or complexity of real EHRs, and should not be used to evaluate clinical model performance or train production models. Its value lies in enabling research on integrating background information into clinical information extraction, within a simplified setting with known and controllable ground-truth structure.
The remainder of this work is structured as follows. Section 2 describes how we generated SimSUM by simulating 10,000 patient encounters by sampling from an expert-defined BN and prompting a large language model. Section 3 analyses the utility of our dataset, and discusses its intended use. Finally, Section 4 states our main conclusions and highlights SimSUM’s limitations.
2 Construction and content
Figure 2: Overview of the full data generating process for the SimSUM dataset. First, the tabular portion of the artificial patient record is sampled from a Bayesian network, where both the structure and the conditional probability distributions were defined by an expert. Afterwards, we construct a prompt describing the symptoms experienced by the patient, as well as their underlying health conditions (but no diagnoses). We ask the GPT-4o large language model to generate a clinical note describing this patient encounter. Finally, we ask to generate a more challenging compact version of the note, mimicking the complexity of real clinical notes by prompting the use of abbreviations and shortcuts. We generate $10{,}000$ of these artificial patient records in total.
Our general methodology for generating the artificial patient records is shown in Fig. 2. We now zoom in on the two major parts of this data generating process. First, Section 2.1 describes how we generated the structured tabular variables through an expert-defined Bayesian network. Then, Section 2.2 dives into how the clinical notes were generated by the large language model. Finally, Section 2.3 describes how we enhanced our dataset by automatically annotating each note with all spans mentioning each symptom.
@@ 2.1 Modeling structured tabular variables with a Bayesian network
Table 1: Description of all 16 tabular variables in our dataset, including their type (underlying condition, diagnosis, symptom, non-clinical, treatment or outcome), and their possible values.
Name Type Description Values
Asthma underlying condition Chronic lung disease in which the airways narrow and swell. yes/no
Smoking underlying condition Whether the patient is a regular smoker of tobacco. yes/no
COPD underlying condition Chronic Obstructive Pulmonary Disease, where airflow from the lungs is obstructed. yes/no
Hay fever underlying condition Allergic rhinitis, irritation of the nose caused by an allergen (e.g. pollen). yes/no
Season non-clinical Season of the year. winter/summer
Pneumonia diagnosis Infection that inflames the air sacs in one or both lungs. yes/no
Common cold diagnosis Upper respiratory tract infection, irritation and swelling of the upper airways. yes/no
Dyspnea symptom Shortness of breath, the feeling of not getting enough air. yes/no
Cough symptom Any type of cough. yes/no
Pain symptom Pain related to the airways or chest area. yes/no
Fever symptom Elevation of body temperature. high/low/none
Nasal symptom Nasal symptoms, such as runny nose or sneezing. yes/no
Antibiotics treatment Whether any type of antibiotics are prescribed to the patient. yes/no
# Days at home outcome How many days the patient ends up staying home as a result of their symptoms and treatment. discrete $(0-...)$
@@
Causal structure We asked an expert to define a Directed Acyclic Graph (DAG) which (partially) models the domain of respiratory diseases in primary care, shown in Fig. 2. In this DAG, a directed arrow between two variables models a causal relation between them. Central to the model are the diagnoses of pneumonia and common cold, which may give rise to five symptoms (dyspnea, cough, pain, fever and nasal). The expert also modeled some relevant underlying conditions which may render a patient more predisposed to certain diagnoses or symptoms: asthma, smoking, COPD and hay fever. Based on the symptoms experienced by a patient, a primary care doctor decides whether to prescribe antibiotics or not. The presence and severity of the symptoms, as well as the prescription of antibiotics as a treatment, influence the outcome, which is the total number of days (# days) that the patient eventually stays home as a result of illness. Finally, there are some non-clinical variables which exert an external influence on the diagnoses, treatment and outcome (season, policy and self-employed). Table 1 summarizes all variables and their meaning, as well as their possible values. While this model does not completely describe the real world, we do believe it to be sufficiently realistic for the purpose of generating a useful artificial dataset.
@@
Probability distribution  To define a data generating mechanism from which we can sample simulated patients, we turn the DAG from Fig. 2 into a Bayesian network by defining a joint probability distribution. This distribution, shown in Equation (1), factorizes into $16$ conditional distributions, one for each variable.
$&\mathcal{P}_{\text{{joint}}}(\text{{asthma, smoking, ..., antibio, \# days}})=\\ &\mathcal{P}(\text{{asthma}})\mathcal{P}(\text{{smoking}})\mathcal{P}(\text{{COPD}}\mid\text{{smoking}})\mathcal{P}(\text{{hayf}})\mathcal{P}(\text{{season}})\mathcal{P}(\text{{cold}}\mid\text{{season}})\\ &\mathcal{P}(\text{{pneu}}\mid\text{{asthma, COPD, season}})\mathcal{P}(\text{{dysp}}\mid\text{{asthma, smoking, COPD, pneu, hayf}})\\ &\mathcal{P}(\text{{cough}}\mid\text{{asthma, smoking, COPD, pneu, cold}})\mathcal{P}(\text{{pain}}\mid\text{{cough, pneu, COPD, cold}})\\ &\mathcal{P}(\text{{fever}}\mid\text{{pneu, cold}})\mathcal{P}(\text{{nasal}}\mid\text{{cold, hayf}})\mathcal{P}(\text{{policy}})\mathcal{P}(\text{{self{\text{-}}empl}})\\ &\mathcal{P}(\text{{antibio}}\mid\text{{policy, dysp, cough, pain, fever}})\\ &\mathcal{P}(\text{{\#days}}\mid\text{{antibio, dysp, cough, pain, fever, nasal, self{\text{-}}empl}})\\ \end{aligned}$
We use four different approaches to parameterize these conditional distributions.
1. Conditional probability table When the variable is discrete and has a limited number of parents, we ask the expert to define a conditional probability table (CPT). Each entry of the table contains the probability for a particular value of the variable, conditional on the combination of values of the parent variables. If the variable has no parents, we just define a prior probability. The probabilities in the tables were filled in based on the expert’s own experience, as well as demographics in their local general practice and Belgium. While we do not expect these probabilities to generalize to the global patient population as a whole, a realistic-looking distribution suffices for our use-case. We provide these tables for the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever, policy and self-employed in Fig. 3.
2. Noisy-OR distribution  For categorical variables with many parents, it becomes infeasible to manually fill in the CPT in a clinically meaningful way, because of the large number of possible combinations of parent values. This is the case for the symptoms dyspnea, cough, pain and nasal in our BN. To circumvent this problem, we define a Noisy-OR distribution, which assumes an independent causal mechanism behind the activation of a symptom through any of its parents [12]. This is a reasonable assumption to make in the case of symptoms with multiple possible causes (parents in the BN): a symptom arises in a patient if any of its possible causes succeeds in activating the symptom through its own independent mechanism. As shown in Equation (2), the parameterization of the Noisy-OR distribution rests on choosing each parameter $p_{i}$, which is the probability that a possible cause $X_{i}$ activates symptom $Y$. As a special case, $p_{0}$, also known as the leak probability, is the probability that symptom $Y$ is activated as the result of another unmodeled cause (outside of all $X_{i}$’s). Note that $x_{i}$ is $1$ when the cause $X_{i}$ is present in the patient (“yes” in our BN), and 00 if not.
$\begin{split}&\mathcal{P}(Y=1\mid X_{1}=x_{1},\dots,X_{k}=x_{k})=1-\mathcal{P}(Y=0\mid X_{1}=x_{1},\dots.,X_{k}=X_{k})\\ &=1-(1-p_{0})(1-p_{1})^{x_{1}}\dots(1-p_{k})^{x_{k}}=\operatorname{Noisy-OR}(p_{0},p_{1},\dots,p_{k};x_{1},\dots,x_{k})\\ &\text{with~{}}x_{1},\dots,x_{k}\in\{0,1\}\end{split}$ (2)
Equations (LABEL:eq:noisy_OR_dysp) through (LABEL:eq:noisy_OR_nasal) define such a Noisy-OR distribution for the symptoms dyspnea, cough, pain and nasal. Note that the symptom fever is fully defined through a CPT, since the expert was able to provide intuition on all possible combinations of its two parent values, eliminating the need for a Noisy-OR distribution.
[…]
$\begin{split}&\mathcal{P}(\text{{pain}}\mid\text{{COPD}},\text{{cough}},\text{{pneu}},\text{{cold}})=\operatorname{Noisy-OR}(p_{0}=0.05,p_{\text{{COPD}}}=0.15,\\ &p_{\text{{cough}}}=0.2,p_{\text{{pneu}}}=0.3,p_{\text{{cold}}}=0.1)\end{split}$ (5)
$\begin{split}&\mathcal{P}(\text{{nasal}}\mid\text{{hayf, cold}})=NoisyOR(p_{0}=0.1,p_{\text{{hayf}}}=0.85,p_{\text{{cold}}}=0.7)\end{split}$ (6)
3. Logistic regression  We model the prescription of antibiotics by mimicking the way the clinician’s suspicion of pneumonia (which needs to be treated with antibiotics) rises when a higher number of symptoms is present in the patient, with some symptoms weighing more than others. Once their level of suspicion reaches a certain threshold, they decide to prescribe treatment. As shown in Equation (7), this process can be modeled using a logistic regression model taking as an input the symptoms dyspnea, cough, pain and fever, as well as the variable policy. Here, $x_{po}$ (policy) can take on the values 1 (high) or 0 (low), $x_{d}$ (dyspnea), $x_{c}$ (cough) and $x_{pa}$ (pain) can take on the value 1 (yes) or 0 (no), and $x_{f}$ (fever) can be 2 (high), 1 (low) or 0 (none).
$\begin{split}&\mathcal{P}(\text{{antibio}}=\text{{yes}}\mid\text{{policy}}=x_{po},\text{{dysp}}=x_{d},\text{{cough}}=x_{c},\text{{pain}}=x_{pa},\text{{fever}}=x_{f})\\ &=\text{Sigmoid}(-3+1\times x_{po}+0.8\times x_{d}+0.665\times x_{c}+0.665\times x_{pa}\\ &+0.9\times(x_{f}==1)+2.25\times(x_{f}==2))\\ &\text{with }x_{po},x_{d},x_{c},x_{pa}\in\{0,1\}\text{, and }x_{f}\in\{0,1,2\}\end{split}$ (7)
We decided on the weights in this model with help of the expert. First, we set the bias term and the coefficient for policy based on the following constraint: if there’s no symptoms at all, and policy is low, then the probability of prescribing antibiotics (due to some other unmodeled cause) should be around 5% . If policy is high, it should be around 10%. All other coefficients were then chosen by the expert based on the relative importance of the symptoms when deciding to prescribe antibiotics. As a final sanity-check, we asked the expert to label a set of test cases with whether or not they would prescribe antibiotics, and compared these with the probability predicted by the model. Table 7 in Appendix A shows these results. We see that the predictions made by the model mostly correspond well with the clinician’s intuition, confirming that the proposed coefficients make sense.
4. Poisson regression  In a similar fashion, we model the # days the patient ends up staying home due to their complaints. Since the outcome is discrete, with most patients staying home for a low number of days, we use a Poisson regression. Assuming a non-linear effect of prescribing antibiotics, we define two separate models: one where antibiotics were not prescribed (Equation (8)), and one where they were (Equation (9)). Both models take as an input the symptoms dysp, cough, pain, nasal and fever, as well as the variable self-employed, and predict a mean number of days $\lambda$, which parameterizes the Poisson distribution.
$\begin{split}&\mathcal{P}(\text{{\# days}}\mid\text{{dysp}}=x_{d},\text{{cough}}=x_{c},\text{{pain}}=x_{pa},\text{{nasal}}=x_{n},\text{{fever}}=x_{f},\\ &\text{{self{\text{-}}empl}}=x_{se},\text{{antibio}}=\text{{no}})=Poisson(\lambda_{0})\\ &\lambda_{0}=exp(0.010+0.64\times x_{d}+0.35\times x_{c}+0.47\times x_{pa}+0.011\times x_{n}\\ &+0.81\times(x_{f}==1)+1.23\times(x_{f}==2)-0.5\times x_{se})\\ &\text{with }x_{d},x_{c},x_{pa},x_{n},x_{se}\in\{0,1\}\text{, and }x_{f}\in\{0,1,2\}\end{split}$ (8)
$\begin{split}&\mathcal{P}(\text{{\# days}}\mid\text{{dysp}}=x_{d},\text{{cough}}=x_{c},\text{{pain}}=x_{pa},\text{{nasal}}=x_{n},\text{{fever}}=x_{f},\\ &\text{{selfempl}}=x_{se},\text{{antibio}}=\text{{yes}})=Poisson(\lambda_{1})\\ &\lambda_{1}=exp(0.16+0.51\times x_{d}+0.42\times x_{c}+0.26\times x_{pa}+0.0051\times x_{n}\\ &+0.24\times(x_{f}==1)+0.57\times(x_{f}==2)-0.5\times x_{se})\\ &\text{with }x_{d},x_{c},x_{pa},x_{n},x_{se}\in\{0,1\}\text{, and }x_{f}\in\{0,1,2\}\end{split}$ (9)
The coefficients for each model were tuned using gradient descent based on the train cases shown in Table 8 in Appendix A. Like before, the expert was asked to (loosely) label these cases with how long they suspected the patient to stay home as a result of these symptoms. The coefficient for the variable self-employed was tuned manually, based on the assumption that being self-employed would shave some days off the predicted number, regardless of the particular symptoms experienced by the patient. As a sanity check, we compared the mean number of days predicted by the model (parameter $\lambda$ in the Poisson model) with the number of days estimated by the expert for a small set of test cases which were not seen during training. The results are shown in Table 9 in Appendix A.
Sampling  The joint probability distribution from Equation (1) is now fully specified. We can use this Bayesian network to randomly sample the tabular portion of a patient record top-down, starting from the root variables without parents at the top and continuing further down. Each value is sampled conditionally on the variable’s parents’ values, using the conditional distributions we have defined. We repeat this process $10{,}000$ times, leaving us with $10{,}000$ artificial patient records consisting of 16 tabular features.